# On the isotopies of tangles in periodic 3-manifolds using finite covers

### v1

@@ 5. Minimal motifs of periodic tangles
Suppose that $X$ is $S^{1}\times D^{2}$, $T^{2}\times I$, or $T^{3}$. Recall the following notation. For links $L_{0},L_{1}\subset X$, we say that $P\colon(X,L_{0})\to(X,L_{1})$ (or $P\colon L_{0}\to L_{1}$ more simply) is a finite covering map if $P\colon X\to X$ is a finite covering map in the usual sense and $L_{0}=P^{-1}(L_{1})$. The universal cover $L_{\infty}$ of a link $L$ in $X$ is called a (singly, doubly, or triply) periodic tangle. Let $L_{0,\infty}$ and $L_{1,\infty}$ be the universal covers of links $L_{0}$ and $L_{1}$ in $X$, respectively. The periodic tangles $L_{0,\infty}$ and $L_{1,\infty}$ are equivalent if there are finite covering maps $P_{0},P_{1}\colon X\to X$ such that there is an admissible homeomorphism $(X,P_{0}^{-1}(L_{0}))\to(X,P_{1}^{-1}(L_{1}))$. Here a homeomorphism $h\colon X\to X$ is admissible if $h$ preserves the orientation in any case, the longitude up to isotopy in the case $X=S^{1}\times D^{2}$, and the boundary components (i.e. the front and back) in the case $X=T^{2}\times I$. We consider the equivalence classes of periodic tangles. A motif of a periodic tangle $L_{\infty}$ is a link in $X$ whose lift in the universal cover of $X$ is equivalent to $L_{\infty}$. A motif $L$ of a periodic tangle $L_{\infty}$ is minimal if there are no smaller motifs. More precisely, $L$ is minimal if the existence of a finite covering map $L\to L^{\prime}$ implies the existence of an admissible homeomorphism $(X,L)\to(X,L^{\prime})$. We show the uniqueness of a minimal motif of a periodic tangle represented by a non-split link. In the case $X=T^{2}\times I$, this is equivalent to the existence of a maximally periodic diagram of a doubly periodic tangle, which is obtained from a diagram on $T^{2}$ of a minimal motif.
@@
Suppose that a periodic tangle $L_{\infty}$ has a motif that is a non-split link in $X$ (we say that $L_{\infty}$ is non-split). Then $L_{\infty}$ has a motif $L$ such that for any motif $L^{\prime}$ of $L_{\infty}$, there is a finite covering map $L^{\prime}\to L$. In particular, $L_{\infty}$ has a minimal motif unique up to homeomorphism.
It is necessary to consider admissible homeomorphisms for the equivalence of singly periodic tangles. For instance, Figure 6 shows two distinct links $L_{0}$ and $L_{1}$ in $S^{1}\times D^{2}$ such that there is a non-admissible homeomorphism between their double covers. However, $L_{0}$ and $L_{1}$ are not non-trivial covers of any other link in $S^{1}\times D^{2}$, since each of them has exactly one null-homotopic component.
There is a split doubly periodic tangle which does not satisfy the above property. For instance, the links $L_{0}$ and $L_{1}$ shown in Figure 8 are motifs of a common doubly periodic tangle. Note that a split summand (a trefoil in the figure) can be freely moved. However, $L_{0}$ and $L_{1}$ have no smaller motifs due to the split summand. The link $L_{0}$ and $L_{1}$ are distinguished by the numbers of components.
@@
Let $L$ be a link in $X$ whose complement is not Seifert fibered. Then there is an upper bound of the degrees of finite covering maps $L\to L^{\prime}$ which depends only on $L$. In particular, there is not a finite covering map $L\to L$ of degree at least two.
@@
We define $\operatorname{vol}(L)$ to be the sum of the volumes of hyperbolic JSJ pieces of the complement of $L$. If the degree of a finite covering map $L\to L^{\prime}$ is $n$, then $\operatorname{vol}(L)=n\operatorname{vol}(L^{\prime})$. Moreover, the set of the volumes of hyperbolic 3-manifolds is well-ordered [24]. In particular, there is a hyperbolic 3-manifold of smallest volume. In fact, Cao and Meyerhoff [8] showed that the smallest volume of orientable cusped hyperbolic 3-manifolds is $2.022v_{\mathrm{tet}}=$ 2.02 …, where $v_{\mathrm{tet}}$ is the volume of a regular ideal tetrahedron. This volume is realized by the figure-eight knot complement. Consequently, if $\operatorname{vol}(L)\neq 0$, the degree of a finite covering map $L\to L^{\prime}$ is at most $\operatorname{vol}(L)/2v_{\mathrm{tet}}$. Hence we may assume that $\operatorname{vol}(L)=0$.
@@
Suppose that every JSJ torus bounds a solid torus. Lemma 2.11 implies that a finite covering map from $L$ can be restricted to a finite covering map from the union of the outermost and adjacent JSJ pieces of $X\setminus\!\!\setminus L$. Hence we may assume that every JSJ tori is a boundary component of the outermost JSJ piece. The outermost JSJ piece is the complement of a link $T_{2}(p,q)$. Consider an adjacent JSJ piece, which is the complement of a link $T_{l}(p^{\prime},q^{\prime})$ in the solid torus. The minimality of JSJ tori implies that $q^{\prime}\neq 0$. The proof of Lemma 4.6 implies that the degree of a finite covering map from $(S^{1}\times D^{2},T_{l}(p^{\prime},q^{\prime}))$ preserving the longitude is at most $q^{\prime}$. Hence the degree of a finite covering map from $L$ is at most $\gcd(p,q)q^{\prime}$.
@@
Suppose that a periodic tangle $L_{\infty}$ is non-split. Let $L_{0}$ and $L_{1}$ be motifs of $L_{\infty}$. Then there is a motif $L$ of $L_{\infty}$ such that there are finite covering maps $L_{0}\to L$ and $L_{1}\to L$.
@@
[…]
Otherwise there is a minimal motif $L$ of $L_{\infty}$ by Lemma 5.2. For any motif $L^{\prime}$ of $L_{\infty}$, there is a motif $L^{\prime\prime}$ such that there are finite covering maps $L\to L^{\prime\prime}$ and $L^{\prime}\to L^{\prime\prime}$ by Lemma 5.3. Since $L$ is minimal, there is a homeomorphism $L^{\prime\prime}\to L$. Hence there is a finite covering map $L^{\prime}\to L$. If another link $L^{\prime}$ is also a minimal motif of $L_{\infty}$, then there are finite covering maps $P\colon L\to L^{\prime}$ and $Q\colon L^{\prime}\to L$. The covering map $Q\circ P\colon L\to L$ is a homeomorphism by Lemma 5.2. Hence $P$ is also a homeomorphism. Therefore a minimal motif of $L_{\infty}$ is unique up to homeomorphism. ∎
@@
Let $P\colon\widetilde{M}\to M$ denote the covering map that is the restriction of the universal covering map $\mathbb{R}^{2}\times I\to T^{2}\times I$. The isometry $g\in\operatorname{Isom}(M)$ lifts an isometry $\tilde{g}\in\operatorname{Isom}(\widetilde{M})$ satisfying that $P\circ\tilde{g}=g\circ P$. Then $\tilde{g}$ acts on a component of $\mathbb{R}^{2}\times\partial I$ by a non-trivial translation. Assume that $g$ has a fixed point in $M$. Then we can take $\tilde{g}$ which has also a fixed point $x$ in $\widetilde{M}$. Take a point $y$ in a horocusp neighborhood of a component of $\mathbb{R}^{2}\times\partial I$. Then the distances between $x=\tilde{g}^{n}(x)$ and $\tilde{g}^{n}(y)$ diverge to infinity when $n\to\infty$. This contradicts the fact that $\tilde{g}$ is an isometry. Hence $g$ has no fixed points in $M$. ∎
@@
Let $M$ be the complement of a hyperbolic link $L$ in $S^{1}\times D^{2}$. Suppose that $g\in\operatorname{Isom}(M)$ is extended to an admissible homeomorphism of $S^{1}\times D^{2}$. Moreover, suppose that the isometry $g$ acts on $S^{1}\times\partial D^{2}$ by a non-trivial translation which does not fix a meridian {}×$\{\ast\}\times\partial D^{2}$ setwise. Then $g$ has no fixed points in $M$.
@@
We next prove in the case $X=T^{3}$. The isometric action of the finite group $G$ on $M$ is extended to a smooth action of $T^{3}$. As in the proof of Theorem 4.4, the geometrization implies that there is a Euclidean metric on $T^{3}$ such that the action of $G$ on $T^{3}$ is isometric. Since the groups $G_{0}$ and $G_{1}$ act on $T^{3}$ by translations, so does $G$. In particular, the action of $G$ on $T^{3}$ is free. Hence the manifold $M/G$ is the complement of a link $L$ in $T^{3}$. Since $G_{i}$ is a subgroup of $G$, the covering map $M/G_{i}\to M/G$ induces a covering map $L_{i}\to L$. ∎
@@
Let $\widetilde{L}$, $L_{0}$, and $L_{1}$ be links in $X$ whose complements are Seifert fibered. Suppose that there are finite covering maps $P_{i}\colon\widetilde{L}\to L_{i}$ for $i=0,1$. If $X=S^{1}\times D^{2}$, suppose that each $P_{i}$ maps a longitude of $X$ to a longitude. Let $G_{i}$ denote the deck transformation groups for the coverings $P_{i}$. Then there are a link $L$ in $X$ and finite covering maps $P^{\prime}_{i}\colon\widetilde{L}\to L_{i}$ and $Q_{i}\colon L_{i}\to L$ that satisfy the following:
@@
    If two components $K$ and $K^{\prime}$ of $\widetilde{L}$ are projected to a common component of $L$ by $P$, then there are components $K_{0},\dots,K_{m}$ of $\widetilde{L}$ such that $K_{j}$ and $K_{j+1}$ are projected to a common component of $L_{i}$ by $P_{i}$ for $i=0$ or $1$.
@@
We prove in the case $X=T^{2}\times I$. In the cases $X=S^{1}\times D^{2}$ and $T^{3}$, we can prove in the same manner. The links $\widetilde{L},L_{0},L_{1}$ are respectively isotopic to $T_{2}(\tilde{p},\tilde{q}),T_{2}(p_{0},q_{0}),T_{2}(p_{1},q_{1})$ by Proposition 3.9. Let $\tilde{d}=\gcd(\tilde{p},\tilde{q})$, $d_{0}=\gcd(p_{0},q_{0})$, and $d_{1}=\gcd(p_{1},q_{1})$. We put $T_{2}(\tilde{p},\tilde{q})$, $T_{2}(p_{0},q_{0})$, and $T_{2}(p_{1},q_{1})$ on the standard positions of the definition in Section 3. For each $i=0,1$, there is $G^{\prime}_{i}$ isomorphic to $G_{i}$ such that $G^{\prime}$ acts on $T^{2}\times I$ by the translations and the actions of $G_{i}$ and $G^{\prime}_{i}$ coincide on the boundary $T^{2}\times\partial I$. Let $G$ be the group generated by the elements of $G^{\prime}_{0}\cup G^{\prime}_{1}$ as translations. Then $G$ is the deck transformation group for a covering map $P\colon(T^{2}\times I,T_{2}(\tilde{p},\tilde{q}))\to(T^{2}\times I,T_{2}(p,q))$, via the covering maps $P^{\prime}_{i}\colon(T^{2}\times I,T_{2}(\tilde{p},\tilde{q}))\to(T^{2}\times I,T_{2}(p_{i},q_{i}))$.
We need to permute the components of $\widetilde{L}$ for Assertion (5). In the above case, the number $d=\gcd(p,q)$ of components of $L=T_{2}(p,q)$ satisfies $d=\gcd(d_{0},d_{1})$. The orbit of each component of $\widetilde{L}$ by the action of $G$ consists of $\tilde{d}/d$ components. We divide the $\tilde{d}$ components of $\widetilde{L}$ maximally into the classes so that two components belong to a common class if they are projected to a common component of $L_{i}$ by $P_{i}$ for $i=0$ or $1$. Two components of different classes should not be projected to a common component by $P$. Each class $C$ is preserved by the action of $G^{\prime}_{i}$. Moreover, each orbit of the action of $G^{\prime}_{i}$ consists of $\tilde{d}/d_{i}$ components. Hence the number $|C|$ of elements is a multiple of $\tilde{d}/d=\operatorname{lcm}(\tilde{d}/d_{0},\tilde{d}/d_{1})$. We can permute the components of $\widetilde{L}$ so that each class is preserved by the action of $G$. Thus we obtain required covering maps $P^{\prime}_{i}$. ∎
@@
Since $L_{0}$ and $L_{1}$ are motifs of $L_{\infty}$, there are a link $\widetilde{L}$ in $X$ and finite covering maps $P_{i}\colon\widetilde{L}\to L_{i}$ for $i=0,1$. The link $\widetilde{L}$ is non-split. We show that there are a link $L$ in $X$ and finite covering maps $P^{\prime}_{i}\colon\widetilde{L}\to L_{i}$ and $Q_{i}\colon L_{i}\to L$ that satisfy Assertions (1)–(4) in Lemma 5.7.
Consider the JSJ decomposition of $X\setminus\!\!\setminus\widetilde{L}$. The deck transformations are restricted to the JSJ pieces. Moreover, they preserve the longitudes on JSJ tori which bound solid tori. The assertion for a hyperbolic piece follows from Lemma 5.6 and the Mostow rigidity. The assertion for a Seifert fibered piece follows from Lemma 5.7. Note that the assertion also holds in the case that the deck transformation groups act on a link component in a disjoint union of solid tori. By combining the actions on the JSJ pieces, we obtain a desired deck transformation on $X\setminus\!\!\setminus\widetilde{L}$. Since knotted hole balls are lifted disjointly by covering maps, the actions are compatible on the boundary of knotted hole balls. Assertion (2) implies that the actions are compatible on a JSJ torus which is layering or bounds a solid torus. The compatibility for pieces adjacent to a Seifert fibered piece requires Assertion (5) in Lemma 5.7. ∎

### v2

@@ 5. Minimal motifs of periodic tangles
Suppose that $X$ is $S^{1}\times D^{2}$, $T^{2}\times I$, or $T^{3}$. Recall the following notation. For links $L_{0},L_{1}\subset X$, we say that $P\colon(X,L_{0})\to(X,L_{1})$ (or $P\colon L_{0}\to L_{1}$ more simply) is a finite covering map if $P\colon X\to X$ is a finite covering map in the usual sense and $L_{0}=P^{-1}(L_{1})$. The universal cover $L_{\infty}$ of a link $L$ in $X$ is called a (singly, doubly, or triply) periodic tangle. Let $L_{0,\infty}$ and $L_{1,\infty}$ be the universal covers of links $L_{0}$ and $L_{1}$ in $X$, respectively. The periodic tangles $L_{0,\infty}$ and $L_{1,\infty}$ are equivalent if there are finite covering maps $P_{0},P_{1}\colon X\to X$ such that there is an admissible homeomorphism $(X,P_{0}^{-1}(L_{0}))\to(X,P_{1}^{-1}(L_{1}))$. Here a homeomorphism $h\colon X\to X$ is admissible if $h$ preserves the orientation in any case, the longitude up to isotopy in the case $X=S^{1}\times D^{2}$, and the boundary components (i.e. the front and back) in the case $X=T^{2}\times I$. We consider the equivalence classes of periodic tangles. A motif of a periodic tangle $L_{\infty}$ is a link in $X$ whose lift in the universal cover of $X$ is equivalent to $L_{\infty}$. A motif $L$ of a periodic tangle $L_{\infty}$ is minimal if there are no smaller motifs. More precisely, $L$ is minimal if the existence of a finite covering map $L\to L^{\prime}$ implies the existence of an admissible homeomorphism $(X,L)\to(X,L^{\prime})$. We show the uniqueness of a minimal motif of a periodic tangle with some conditions. In the case $X=T^{2}\times I$, this is equivalent to the existence of a maximally periodic diagram of a doubly periodic tangle, which is obtained from a diagram on $T^{2}$ of a minimal motif.
@@
Let $L_{\infty}$ be a non-split periodic tangle with a motif $L\subset X=S^{1}\times D^{2}$, $T^{2}\times I$, or $T^{3}$. Suppose that each Seifert fibered JSJ piece $M$ of the complement of $L$ such that the homomorphism $\iota_{*}\colon H_{1}(M,\mathbb{Z})\to H_{1}(X,\mathbb{Z})$ induced by the inclusion map is non-trivial can be re-embedded as the complement of $T_{0}(p,0)$, $T_{1}(0,q)$, $T_{2}(p,q)$, or $T_{3}(p,q,r)$. Then $L_{\infty}$ has a motif $L_{\min}$ such that for any motif $L^{\prime}$ of $L_{\infty}$, there is a finite covering map $L^{\prime}\to L_{\min}$. In particular, $L_{\infty}$ has a minimal motif unique up to homeomorphism.
We need to consider admissible homeomorphisms for the equivalence of singly periodic tangles. For instance, Figure 6 shows two distinct links $L_{0}$ and $L_{1}$ in $S^{1}\times D^{2}$ such that there is a non-admissible homeomorphism between their double covers. However, $L_{0}$ and $L_{1}$ are not non-trivial covers of any other link in $S^{1}\times D^{2}$, since each of them has exactly one null-homotopic component.
We need the condition for splitness and JSJ pieces to avoid the following counterexamples. However, some periodic tangles that do not satisfy the assumption of Theorem 5.1 may accidentally have unique minimal motifs. A necessary and sufficient condition for having a unique minimal motif would be too complicated.
There is a split doubly periodic tangle whose minimal motifs are not unique. For instance, the links $L_{0}$ and $L_{1}$ shown in Figure 8 are motifs of a common doubly periodic tangle. Note that a split summand (a trefoil in the figure) can be freely moved. However, $L_{0}$ and $L_{1}$ have no smaller motifs due to the split summand. The links $L_{0}$ and $L_{1}$ are distinguished by the number of components. We remark that splitness does not matter for singly periodic tangles.
@@
Moreover, the links $T_{0}(3,1)$ and $T_{1}(2,1)$ in $S^{1}\times D^{2}$ are distinguished by the number of components and minimal by the classification by Lemma 3.1 and Proposition 3.2. Since the triple cover $T_{0}(3,3)$ of $T_{0}(3,1)$ is isotopic to the double cover $T_{1}(2,2)$ of $T_{1}(2,1)$, the links $T_{0}(3,1)$ and $T_{1}(2,1)$ are motifs of a common singly periodic tangle. This corresponds to the relation $(\sigma_{1}\sigma_{2})^{3}=(\sigma_{1}\sigma_{2}\sigma_{1})^{2}$ in the braid group $B_{3}$. More generally, the $p$-sheeted cover $T_{0}(p,p)$ of $T_{0}(p,1)$ is isotopic to the p1p-1-sheeted cover $T_{1}(p-1,p-1)$ of $T_{1}(p-1,1)$ for $p\geq 2$. There does not exist such coincidence for $T_{0}(p,0)$ and $T_{1}(0,q)$. If a periodic tangle $L_{\infty}$ has a motif that is a connected sum of a hyperbolic link in $X$ and hyperbolic knots in $S^{3}$, then each Seifert fibered JSJ piece is $T_{1}(0,q)$, and so $L_{\infty}$ has a unique minimal motif by Theorem 5.1.
@@
Let $L$ be a link in $X$ whose complement is not Seifert fibered. Then there is an upper bound of the degrees of finite covering maps $L\to L^{\prime}$ which depends only on $L$. In particular, there is no finite covering map $L\to L$ of degree at least two.
@@
We define $\operatorname{vol}(L)$ to be the sum of the volumes of hyperbolic JSJ pieces of the complement of $L$. If the degree of a finite covering map $L\to L^{\prime}$ is $n$, then $\operatorname{vol}(L)=n\operatorname{vol}(L^{\prime})$. Moreover, the set of the volumes of hyperbolic 3-manifolds is well-ordered [27]. In particular, there is a hyperbolic 3-manifold of smallest volume. In fact, Cao and Meyerhoff [9] showed that the smallest volume of orientable cusped hyperbolic 3-manifolds is $2v_{\mathrm{tet}}=2.02$..., where $v_{\mathrm{tet}}$ is the volume of a regular ideal tetrahedron. This volume is realized by the figure-eight knot complement. Consequently, if $\operatorname{vol}(L)\neq 0$, the degree of a finite covering map $L\to L^{\prime}$ is at most $\operatorname{vol}(L)/2v_{\mathrm{tet}}$. Hence we may assume that $\operatorname{vol}(L)=0$.
@@
Suppose that every JSJ torus bounds a solid torus. Lemma 2.11 implies that a finite covering map from $L$ can be restricted to a finite covering map from the union of the outermost and adjacent JSJ pieces of $X\setminus\!\!\setminus L$. Hence we may assume that every JSJ torus is a boundary component of the outermost JSJ piece. The outermost JSJ piece is the complement of a link $T_{2}(p,q)$. Consider an adjacent JSJ piece, which is the complement of a link $T_{l}(p^{\prime},q^{\prime})$ in the solid torus. The minimality of JSJ tori implies that $q^{\prime}\neq 0$. The proof of Lemma 4.6 implies that the degree of a finite covering map from $(S^{1}\times D^{2},T_{l}(p^{\prime},q^{\prime}))$ preserving the longitude is at most $q^{\prime}$. Hence the degree of a finite covering map from $L$ is at most $\gcd(p,q)q^{\prime}$.
@@
Let $L_{\infty}$ be a non-split periodic tangle with a motif $L\subset X=S^{1}\times D^{2}$, $T^{2}\times I$, or $T^{3}$. Suppose that each Seifert fibered JSJ piece $M$ of the complement of $L$ such that the homomorphism $\iota_{*}\colon H_{1}(M,\mathbb{Z})\to H_{1}(X,\mathbb{Z})$ induced by the inclusion map is non-trivial can be re-embedded as $T_{0}(p,0)$, $T_{1}(0,q)$, $T_{2}(p,q)$, or $T_{3}(p,q,r)$. Let $L_{0}$ and $L_{1}$ be motifs of $L_{\infty}$. Then there is a motif $\widehat{L}$ of $L_{\infty}$ such that there are finite covering maps $L_{0}\to\widehat{L}$ and $L_{1}\to\widehat{L}$.
@@
If the link $T_{0}(p,0)$, $T_{1}(0,q)$, $T_{2}(p,q)$, or $T_{3}(p,q,r)$ is a motif of $L_{\infty}$, then $T_{0}(1,0)$, $T_{1}(0,1)$, $T_{2}(1,0)$, or $T_{3}(1,0,0)$ is a unique minimal motif of $L_{\infty}$. Otherwise there is a minimal motif $L_{\min}$ of $L_{\infty}$ by Lemma 5.2. For any motif $L^{\prime}$ of $L_{\infty}$, there is a motif $\widehat{L}$ such that there are finite covering maps $L_{\min}\to\widehat{L}$ and $L^{\prime}\to\widehat{L}$ by Lemma 5.3. Since $L_{\min}$ is minimal, there is a homeomorphism $\widehat{L}\to L_{\min}$. Hence there is a finite covering map $L^{\prime}\to L_{\min}$. If $L^{\prime}$ is also a minimal motif of $L_{\infty}$, then there are finite covering maps $P\colon L_{\min}\to L^{\prime}$ and $Q\colon L^{\prime}\to L_{\min}$. The covering map $Q\circ P\colon L_{\min}\to L_{\min}$ is a homeomorphism by Lemma 5.2. Hence $P$ is also a homeomorphism. Therefore a minimal motif of $L_{\infty}$ is unique up to homeomorphism. ∎
@@
[…]
@@
Let $M$ be the complement of a hyperbolic link $L$ in $S^{1}\times D^{2}$. Suppose that $g\in\operatorname{Isom}(M)$ is extended to an admissible homeomorphism of $S^{1}\times D^{2}$. Moreover, suppose that the isometry $g$ acts on $S^{1}\times\partial D^{2}$ by a non-trivial translation which does not fix a meridian $\{\ast\}\times\partial D^{2}$ setwise. Then $g$ has no fixed points in $M$.
@@
We next prove in the case $X=T^{3}$. The isometric action of the finite group $G$ on $M$ is extended to a smooth action of $T^{3}$. As in the proof of Theorem 4.4, the result of Meeks and Scott [49] implies that there is a Euclidean metric on $T^{3}$ such that the action of $G$ on $T^{3}$ is isometric. Since the groups $G_{0}$ and $G_{1}$ act on $T^{3}$ by translations, so does $G$. In particular, the action of $G$ on $T^{3}$ is free. Hence the manifold $M/G$ is the complement of a link $L$ in $T^{3}$. Since $G_{i}$ is a subgroup of $G$, the covering map $M/G_{i}\to M/G$ induces a covering map $L_{i}\to L$. ∎
@@
Let $\widetilde{L}$, $L_{0}$, and $L_{1}$ be links in $X$ whose complements are $T_{0}(p,0)$, $T_{1}(0,q)$, $T_{2}(p,q)$, or $T_{3}(p,q,r)$. Suppose that there are finite covering maps $P_{i}\colon\widetilde{L}\to L_{i}$ for $i=0,1$. If $X=S^{1}\times D^{2}$, suppose that each $P_{i}$ maps a longitude of $X$ to a longitude. Let $G_{i}$ denote the deck transformation groups for the coverings $P_{i}$. Then there are a link $\widehat{L}$ in $X$ and finite covering maps $P^{\prime}_{i}\colon\widetilde{L}\to L_{i}$ and $Q_{i}\colon L_{i}\to\widehat{L}$ that satisfy the following:
@@
    If two components $K$ and $K^{\prime}$ of $\widetilde{L}$ are projected to a common component of $\widehat{L}$ by $P$, then there are components $K_{0},\dots,K_{m}$ of $\widetilde{L}$ such that $K_{j}$ and $K_{j+1}$ are projected to a common component of $L_{i}$ by $P_{i}$ for $i=0$ or $1$.
@@
[…]
We need to permute the components of $\widetilde{L}$ for Assertion (5). In the above case, the number $d=\gcd(p,q)$ of components of $\widehat{L}=T_{2}(p,q)$ satisfies $d=\gcd(d_{0},d_{1})$. The orbit of each component of $\widetilde{L}$ by the action of $G$ consists of $\tilde{d}/d$ components. We divide the $\tilde{d}$ components of $\widetilde{L}$ maximally into the classes so that two components belong to a common class if they are projected to a common component of $L_{i}$ by $P_{i}$ for $i=0$ or $1$. Two components of different classes should not be projected to a common component by $P$. Each class $C$ is preserved by the action of $G^{\prime}_{i}$. Moreover, each orbit of the action of $G^{\prime}_{i}$ consists of $\tilde{d}/d_{i}$ components. Hence the number $|C|$ of elements is a multiple of $\tilde{d}/d=\operatorname{lcm}(\tilde{d}/d_{0},\tilde{d}/d_{1})$. We can permute the components of $\widetilde{L}$ so that each class is preserved by the action of $G$. Thus we obtain the required covering maps $P^{\prime}_{i}$. ∎
@@
Since $L_{0}$ and $L_{1}$ are motifs of $L_{\infty}$, there are a link $\widetilde{L}$ in $X$ and finite covering maps $P_{i}\colon\widetilde{L}\to L_{i}$ for $i=0,1$. The link $\widetilde{L}$ is non-split. We show that there are a link $\widehat{L}$ in $X$ and finite covering maps $P^{\prime}_{i}\colon\widetilde{L}\to L_{i}$ and $Q_{i}\colon L_{i}\to\widehat{L}$ that satisfy Assertions (1)–(4) in Lemma 5.7.
Consider the JSJ decomposition of $X\setminus\!\!\setminus\widetilde{L}$. The deck transformations are restricted to the JSJ pieces. Moreover, they preserve the longitudes on JSJ tori that bound solid tori. The assertion for a hyperbolic piece follows from Lemma 5.6 and the Mostow rigidity theorem. The assertion for a Seifert fibered piece that can be re-embedded as $T_{0}(p,0)$, $T_{1}(0,q)$, $T_{2}(p,q)$, or $T_{3}(p,q,r)$ follows from Lemma 5.7. Note that the assertion also holds in the case that the deck transformation groups act on a link component in a disjoint union of solid tori. By combining the actions on the JSJ pieces, we obtain a desired deck transformation on $X\setminus\!\!\setminus\widetilde{L}$. Since knotted hole balls are lifted disjointly by covering maps, the actions are compatible on the boundary of knotted hole balls. If $\iota_{*}\colon H_{1}(M,\mathbb{Z})\to H_{1}(X,\mathbb{Z})$ is trivial for a Seifert fibered piece $M$ in $X\setminus\!\!\setminus L_{i}$, then $M$ is also lifted disjointly by the covering map. Assertion (2) implies that the actions are compatible on a JSJ torus that is layering or bounds a solid torus. The compatibility for pieces adjacent to a Seifert fibered piece requires Assertion (5) in Lemma 5.7. ∎